# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v2

@@ 3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn projected SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (i.e., they monitor validation loss, gradient norms, or empirical convergence patterns) and their guarantees, when available, typically hold only for a pre-specified time or for a specific deterministic stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies. Throughout this section, we work with the observable confidence sequence $U_{t}^{\mathrm{obs}}(\alpha)$, and Assumption 2.1 can be imposed on the observable variance proxy $\widetilde{\Sigma}_{t}^{2}=\sigma^{2}R_{x}^{2}V_{t}$.
Corollary 2.5 provides an anytime-valid upper confidence sequence $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ in (4). Combining this with the convexity bound $f(\bar{x}_{t})-f^{\star}\leq\bar{F}_{t}$, we obtain, for any fixed $\alpha\in(0,1)$,
@@
Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,1)$, it is therefore natural to stop as soon as the certificate $U_{t}^{\mathrm{obs}}(\alpha)$ falls below $\varepsilon$. As already foreshadowed in the introduction, we formalize this through the stopping time
@@
Suppose Assumptions 1.1, 1.2, and 2.1 (with $\Sigma_{t}^{2}$ replaced by $\widetilde{\Sigma}_{t}^{2}$) hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.5. Fix $\varepsilon>0$ and $\alpha\in(0,0.6)$, and define the stopping time $\tau_{\varepsilon}$ by (31). Then:
@@
We first show that $\tau_{\varepsilon}$ is almost surely finite. By Proposition 2.3, for any fixed $\alpha\in(0,0.6)$, we have
@@
Theorem 3.1 shows that projected SGD equipped with the stopping rule $\tau_{\varepsilon}$ is a data-dependent $\varepsilon$-optimization procedure: one may monitor $U_{t}^{\mathrm{obs}}(\alpha)$ online and stop as soon as it falls below $\varepsilon$, without specifying a horizon in advance or invalidating the $1-\alpha$ guarantee. This yields fully implementable stopping rules based only on running sums of $\eta_{t}$, $\eta_{t}^{2}$, and the observed gradient norms, with the expected tradeoff between adaptivity and reliance on global parameters.
In the following subsection we study the stopping-time complexity associated with the rule $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$. We present explicit bounds with deterministic constants via a strengthening of our assumptions on the stochastic gradients.
@@ 3.1 Stopping-time complexity
Let Assumptions 3.2, 1.2, and 2.1 (with $\Sigma_{t}^{2}$ replaced by $\widetilde{\Sigma}_{t}^{2}$) hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.5. Moreover, define
$K_{\alpha}:=\frac{1}{2}\Bigg(C\sqrt{\widetilde{\Sigma}_{\infty}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log\!\big(\mathrm{e}+\widetilde{\Sigma}_{\infty}^{2}\big)\Big)}+R_{x}^{2}+G^{2}V_{\infty}\Bigg),$ (33)
with constant $C$ defined in (10), $\widetilde{\Sigma}_{\infty}^{2}:=4G^{2}R_{x}^{2}V_{\infty}$, and $V_{\infty}:=\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Then, for every target accuracy $\varepsilon>0$,
@@ Appendix A Technical lemmas
Let $\widetilde{\Sigma}_{\infty}^{2}>1$ and define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\}.$
Fix any $L_{\alpha}\geq 1$. Then, for all $v$ with $L_{\alpha}<v\leq\widetilde{\Sigma}_{\infty}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;c_{1}\,\log\Big(\log(v+e)\Big).$
@@
First note that for $v>L_{\alpha}/e$ we have $v/L_{\alpha}>1/e$, hence $\log(v/L_{\alpha})>-1$, so $\log(v/L_{\alpha})+1>0$ and $\log(\log(v/L_{\alpha})+1)$ is well-defined. Moreover, $v+e>e$ implies $\log(v+e)>1$ and thus $\log(\log(v+e))>0$. We now prove the inequality. Since $L_{\alpha}\geq 1$ and the logarithm is an increasing function,
$L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty}^{2}\ \implies\ \log(v/L_{\alpha})+1\leq\log\widetilde{\Sigma}_{\infty}^{2}+1\ \implies\ \log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty}^{2}+1).$
@@
$v\geq 1\ \implies\ \log(v+e)\geq\log(e+1)>1\ \implies\ \log(\log(v+e))\geq\log(\log(e+1)).$
@@
$c_{1}\;\geq\;\frac{\log(\log\widetilde{\Sigma}_{\infty}^{2}+1)}{\log(\log(e+1))}\ \implies\ \log(\log\widetilde{\Sigma}_{\infty}^{2}+1)\;\leq\;c_{1}\,\log(\log(e+1)).$
Combining the last three displays, we obtain for all $L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty}^{2}+1)\;\leq\;c_{1}\,\log(\log(e+1))\;\leq\;c_{1}\,\log(\log(v+e)),$
@@
Since $L_{\alpha}\geq 1$, we have $\nu_{0}=2(L_{\alpha}+1)\geq 4$. For any $v\geq 4$, we have $e^{v}\geq e+v$, which implies that $\log(e+v)\leq v$. Since $\log(e+v)>1$, and therefore $\log\Big(\log(e+v)\Big)\leq\log v$ for all $v\geq 4$. Thus, for $v\geq\nu_{0}$, we have
$v-L_{\alpha}-\log\Big(\log(e+v)\Big)\geq v-L_{\alpha}-\log v=:g(v)$
@@
Therefore $v-L_{\alpha}-\log(\log(e+v))>0$ for all $v\geq\nu_{0}$, which concludes the proof. ∎

### v3

@@ 3 Anytime-Valid Stopping Rules
A central motivation for our confidence-sequence approach is to turn projected SGD into an anytime-valid optimization procedure: instead of committing to a fixed horizon in advance, the user would like to stop the algorithm based on the information actually observed along the trajectory. In practice, most stopping rules for SGD are heuristic (i.e., they monitor validation loss, gradient norms, or empirical convergence patterns) and their guarantees, when available, typically hold only for a pre-specified time or for a specific deterministic stopping rule. In contrast, a confidence sequence is designed precisely to support optional stopping: it provides a high-probability guarantee that remains valid under arbitrary data-dependent monitoring and stopping strategies.
Corollary 2.4 provides an observable anytime-valid upper confidence sequence $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ in (4). Combining this with the convexity bound $f(\bar{x}_{t})-f^{\star}\leq\bar{F}_{t}$, we obtain, for any fixed $\alpha\in(0,1)$,
@@
Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, it is therefore natural to stop as soon as the certificate $U_{t}^{\mathrm{obs}}(\alpha)$ falls below $\varepsilon$. As already foreshadowed in the introduction, we formalize this through the stopping time
@@
Suppose Assumptions 1.1 and 1.2 hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.4. Fix $\varepsilon>0$ and $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, and define the stopping time $\tau_{\varepsilon}$ by (32). Then:
@@
We first show that $\tau_{\varepsilon}$ is almost surely finite. Following the same reasoning as in Proposition 2.2, for any fixed $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, we have
@@
Theorem 3.1 shows that projected SGD equipped with the stopping rule $\tau_{\varepsilon}$ is a data-dependent $\varepsilon$-optimal procedure: one may monitor $U_{t}^{\mathrm{obs}}(\alpha)$ online and stop as soon as it falls below $\varepsilon$, without specifying a horizon in advance or invalidating the $1-\alpha$ guarantee. This yields fully implementable stopping rules based only on running sums of $\eta_{t}$, $\eta_{t}^{2}$, and the observed gradient norms, with the usual tradeoff between adaptivity and reliance on global problem parameters.
In the next subsection we study the stopping-time complexity associated with the rule $U_{t}^{\mathrm{obs}}(\alpha)\leq\varepsilon$. We present explicit bounds with deterministic constants via a strengthening of our assumptions on the stochastic gradients.
@@ 3.1 Stopping-time complexity
Let Assumptions 3.2 and 1.2 hold, and let $\{U_{t}^{\mathrm{obs}}(\alpha)\}_{t\geq 1}$ be the anytime-valid upper confidence sequence from Corollary 2.4. Moreover, define
$K_{\alpha}:=\frac{1}{2}\Bigg(C\sqrt{\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log\!\big(\mathrm{e}+\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\big)\Big)}+R_{x}^{2}+G^{2}V_{\infty}\Bigg),$ (34)
with constant $C$ defined in (10) and $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}:=\max\{4G^{2}R_{x}^{2}V_{\infty},2(\log\tfrac{2}{\alpha}+1)\}$. Then, for every target accuracy $\varepsilon>0$,
@@ Appendix A Technical lemmas
Let $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}>1$ and define
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\}.$
Fix any $L_{\alpha}\geq 1$. Then, for all $v$ with $L_{\alpha}<v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;c_{1}\,\log\Big(\log(v+\mathrm{e})\Big).$
@@
First note that for $v>L_{\alpha}/\mathrm{e}$ we have $v/L_{\alpha}>1/\mathrm{e}$, hence $\log(v/L_{\alpha})>-1$, so $\log(v/L_{\alpha})+1>0$ and $\log(\log(v/L_{\alpha})+1)$ is well-defined. Moreover, $v+\mathrm{e}>\mathrm{e}$ implies $\log(v+\mathrm{e})>1$ and thus $\log(\log(v+\mathrm{e}))>0$. We now prove the inequality. Since $L_{\alpha}\geq 1$ and the logarithm is an increasing function,
$L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}\ \implies\ \log(v/L_{\alpha})+1\leq\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\ \implies\ \log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1).$
@@
$v\geq 1\ \implies\ \log(v+\mathrm{e})\geq\log(\mathrm{e}+1)>1\ \implies\ \log(\log(v+\mathrm{e}))\geq\log(\log(\mathrm{e}+1)).$
@@
$c_{1}\;\geq\;\frac{\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)}{\log(\log(\mathrm{e}+1))}\ \implies\ \log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)\;\leq\;c_{1}\,\log(\log(\mathrm{e}+1)).$
Combining the last three displays, we obtain for all $L_{\alpha}\leq v\leq\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$,
$\log\Big(\log(v/L_{\alpha})+1\Big)\;\leq\;\log(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1)\;\leq\;c_{1}\,\log(\log(\mathrm{e}+1))\;\leq\;c_{1}\,\log(\log(v+\mathrm{e})),$
@@
Since $L_{\alpha}\geq 1$, we have $\nu_{0}=2(L_{\alpha}+1)\geq 4$. For any $v\geq 4$, we have $\mathrm{e}^{v}\geq\mathrm{e}+v$, which implies that $\log(\mathrm{e}+v)\leq v$. Since $\log(\mathrm{e}+v)>1$, and therefore $\log\Big(\log(\mathrm{e}+v)\Big)\leq\log v$ for all $v\geq 4$. Thus, for $v\geq\nu_{0}$, we have
$v-L_{\alpha}-\log\Big(\log(\mathrm{e}+v)\Big)\geq v-L_{\alpha}-\log v=:g(v)$
@@
Therefore $v-L_{\alpha}-\log(\log(\mathrm{e}+v))>0$ for all $v\geq\nu_{0}$, which concludes the proof. ∎